### v1

[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
Explainable deep learning improves human mental models of self-driving cars
[…]
Summary
Self-driving cars increasingly rely on deep neural networks to achieve human-like driving [phan2023driveirl, marcu2023lingoqa]. However, the opacity of such black-box motion planners makes it challenging for the human behind the wheel to accurately anticipate when they will fail [nikolaidis2012human, major2020expect, paleja2021utility], with potentially catastrophic consequences [Titcomb2019, Fang2023, templeton2024waymo]. Here, we introduce concept-wrapper network (i.e., CW-Net), a method for explaining the behavior of black-box motion planners by grounding their reasoning in human-interpretable concepts. We deploy CW-Net on a real self-driving car and show that the resulting explanations refine the human driver’s mental model of the car, allowing them to better predict its behavior and adjust their own behavior accordingly. Unlike previous work using toy domains or simulations [kenny2023towards, kim2023and, atakishiyev2024explainable], our study presents the first real-world demonstration of how to build authentic autonomous vehicles (AVs) that give interpretable, causally faithful explanations for their decisions, without sacrificing performance. We anticipate our method could be applied to other safety-critical systems with a human in the loop, such as autonomous drones and robotic surgeons. Overall, our study suggests a pathway to explainability for autonomous agents as a whole, which can help make them more transparent, their deployment safer, and their usage more ethical.
keywords:
Self-driving cars, Mental model alignment, Interpretable ML, AI safety
1 Introduction
There are hundreds of companies developing autonomous vehicle (AV) technology globally [badue2021self], promising to revolutionize transportation for everyone. However, the complexity of fully driverless autonomy has prompted an industry shift towards advanced driver-assistance systems, which require successful communication between the AV and human driver. This is made increasingly difficult by the adoption of deep neural networks in AVs for planning and decision making, the core cognitive functions that determine driving behavior  [atakishiyev2024explainable]. Deep learning allows motion planners to learn the nuances of human driving behavior from data, but the implicit nature of the learned driving policies makes it challenging to understand the causes of their decisions and to predict their behavior.
A lack of effective communication between the AV and the human driver has contributed to multiple high-profile incidents, some resulting in fatalities [Titcomb2019, Fang2023, templeton2024waymo], highlighting the urgent need to make deep motion planners interpretable [atakishiyev2024explainable]. Previous studies have sought to address this need using surveys and simulated scenarios [koo2015did, koo2016understanding, wiegand2020d, wang2021human, schneider2021increasing, schneider2021explain, omeiza2021towards, zemni2023octet], a human driver emulating the AV [kim2023and, schneider2023don], or large language models providing post-hoc explanations [marcu2023lingoqa]. However, these studies were theoretical, did not provide faithful explanations of the AV’s reasoning process, or were only evaluated in simulation. This leaves open the question of how to provide understandable and useful explanations for the decisions of deep motion planners deployed in real self-driving cars.
To answer this question, we scale up our recent work on interpretable-by-design deep reinforcement learning with prototype-wrapper networks (PW-Nets) [kenny2023towards], using motifs from the literature on concept-bottleneck models [yuksekgonulpost]. Our key proposal is to ground the reasoning of black-box motion planners in human-interpretable concepts, such as “Approaching stopped vehicle” or “Close to cyclist”. This method is rooted in case-based reasoning, a classical artificial intelligence (AI) approach [leake1996case, keane2019case, sormo2005explanation] inspired by cognitive models of human reasoning and memory [schank1983dynamic]. It results in causal explanations, such as “I chose to stop based on recognizing that we are approaching a stopped vehicle.”. In our tests, these explanations help align a safety driver’s mental model of the AV with its actual internal decision-making process, increasing transparency and predictability. Importantly, this approach can be applied to arbitrary pre-trained deep neural networks, does not require retraining from scratch, and does not degrade performance of the original black-box planner.
We apply our proposed method, CW-Net (short for concept-wrapper network), to a deep motion planner trained to imitate human driving behavior using inverse reinforcement learning [phan2023driveirl]. We replace the final (reward) layer of the pretrained deep neural network with a concept classifier, followed by a new reward layer. We then jointly train the classifier and the new reward layer to predict scenario types and driving decisions, respectively, without modifying the rest of the network. Evaluation on a large-scale benchmark [karnchanachari2024towards] confirms that CW-Net is able to classify concepts without compromising driving behavior. To study the utility of the explanations, we then deploy CW-Net on a real self-driving car [heim2024lab2car]. We demonstrate three situations in which the driver has an inaccurate mental model of the motion planner, which is subsequently corrected by the explanations. This ultimately changes the behavior of the driver, for example, by increasing their vigilance in certain situations. Finally, we confirm the statistical significance of these results in an online study (N=120). Overall, our work demonstrates how explainable AI can help users of advanced autonomous systems better understand their behavior in naturalistic settings [endsley2017autonomous], while also providing insights that can potentially accelerate the development and refinement of such systems.
2 Black-box motion planner
Figure 1: Planner architecture. a. Autonomous vehicle stack. Sensory input is processed by the perception module to generate scene context $s$. The planning module processes $s$ to compute trajectory $\hat{\tau}$, which is followed by the control module. b. Black-box motion planner. $s$ is fed to trajectory generator $G$, which produces candidate trajectories $\{\tau_{1}\dots\tau_{k}\}$, and scene encoder $H$, which produces a scene embedding $\mathbf{h}$. These are fed into encoder $E$, which produces scene-trajectory embeddings $\mathbf{z}_{i}$ for each $\tau_{i}$, which are in turn fed into the reward layer $R$. $R$ computes a reward $r_{i}$ for each trajectory. The model is trained to output higher rewards for trajectories closer to the the ground-truth human trajectories. c. CW-Net. Identical to b., except $\mathbf{z}_{i}$ is fed to a concept classifier $C$, which produces concept assignments $\mathbf{c}_{i}$. These are fed to a new reward layer $R^{\prime}$ to produce rewards $r^{\prime}_{i}$. In parallel, $\mathbf{c}_{i}$ is processed to generate the explanations. The reward layer is trained to prefer the same trajectories as the black-box planner, while the concept layer is supervised with ground-truth scenario labels. The weights of the faded components are frozen during training. CLOSE, “Close to another vehicle”. ASV, “Approaching stopped vehicle”. BIKE, “Close to cyclist”.
We focus on the motion planning module of the AV stack (Figure 1a), which takes as input a scene context $s$ and outputs a trajectory $\hat{\tau}$. $s$ is a symbolic object-oriented representation of the scene computed by the perception module, while $\hat{\tau}$ is the trajectory that the subsequent controller module should follow. We use a deep neural network architecture consisting of a scene encoder $H(s)\rightarrow\mathbf{h}$ and a trajectory generator $G(s)\rightarrow\{\tau_{1}\dots\tau_{k}\}$, followed by a scene-trajectory encoder $E(\mathbf{h},\tau_{i})\rightarrow\mathbf{z}_{i}$ and a final reward layer $R(\mathbf{z}_{i})\rightarrow r_{i}$ (Figure 1b). $H$ computes a scene embedding $\mathbf{h}$ and $G$ computes a set of $k$ candidate trajectories $\{\tau_{1}\dots\tau_{k}\}$. Those are combined in $E$ to compute an embedding $\mathbf{z}_{i}$ for each trajectory $\tau_{i}$. Finally, $R$ computes an estimated reward $r_{i}$ for each trajectory $\tau_{i}$, quantifying how human-like it is. In other words, $r_{i}$ is higher when $\tau_{i}$ is more similar to how a human would drive in this situation.
During inference, the trajectory with highest reward is selected on each iteration:
$\hat{\tau}=\tau_{\hat{i}}\quad\text{such that}\quad\hat{i}=\underset{i\in{1,...,k}}{\arg\max}\,r_{i},\quad r_{i}=R(E(\mathbf{h},\tau_{i}))$
We train the planner on 80 hours of human expert driving using inverse reinforcement learning [phan2023driveirl].
3 Planning over human-friendly concepts
To make the planner more interpretable, we replace $R$ with a concept classifier $C(\mathbf{z}_{i})\rightarrow\mathbf{c}_{i}$, followed by a new reward layer $R^{\prime}(\mathbf{c}_{i})\rightarrow r^{\prime}_{i}$ (Figure 1c). The concept classifier $C$ computes a logit vector $\mathbf{c}_{i}$ which is passed through a softmax and/or sigmoid layer which assigns probabilities to different human-interpretable concepts. Since $R^{\prime}$ computes trajectory rewards from $\mathbf{c}_{i}$, the final decisions are based solely on these concept assignments and hence they constitute a causally faithful explanation. The rest of the network remains the same.
Similarly to the black-box planner, trajectories are selected according to:
$\hat{\tau}=\tau_{\hat{i}}\quad\text{such that}\quad\hat{i}=\underset{i\in{1,...,k}}{\arg\max}\,r^{\prime}_{i},\quad r^{\prime}_{i}=R^{\prime}(C(E(\mathbf{h},\tau_{i})))$
We train CW-Net to jointly predict concept labels and mimic the driving decisions of the black-box planner. Specifically, $\mathbf{c}_{i}$ is supervised with multinomial labels corresponding to types of scenarios, such as “Approaching stopped vehicle” or “Close to cyclist”. This ensures that CW-Net assigns a unique interpretable concept to each unit in $\mathbf{c}_{i}$. At the same time, $r^{\prime}_{i}$ is supervised with trajectories selected by the black-box planner using a cross-entropy loss. During training, the rest of the deep neural network ($H$, $G$, and $E$) is kept frozen (see Section 8.2 for more details).
4 Evaluation in simulation
As a baseline, we first evaluated the black-box planner (without CW-Net) using closed-loop simulations on the nuPlan dataset [karnchanachari2024towards] (Table S1). Overall, the results were competitive with the top submissions to the nuPlan challenge, although performance was slightly lacking when starting from a stop (see Section Simulation Results). This suggests that there is room for improvement and, importantly, opportunities to study explanations of undesirable behavior.
We then evaluated the driving performance of CW-Net wrapped around the black-box planner (Table S1). The results were equivalent, with less than 1% difference across all metrics, confirming our method did not degrade driving performance. We also evaluated concept classification on held-out datasets (Table S2/S3). Mean accuracy was 54%, with 23% precision, 77% recall, and an F1 score of 0.31 (see Section 8.3). Overall, these results indicate that CW-Net can be used to ground the decision making of high-performance deep motion planners in human-interpretable concepts without sacrificing driving performance.
5 Explanations in real-world deployment
Figure 2: Deployment setup. A safety driver, a support engineer, and a researcher were present. The safety driver drove the AV manually between road tests, engaged self-driving mode at the start of each test, monitored AV performance during the test, and took over in case of unsafe driving. The support engineer deployed CW-Net and set scenario destinations. The researcher directed testing. The dashboard included a map with overlaid object detections ($s$) from the perception module and the output trajectory ($\hat{\tau}$). Explanations $\mathbf{c}_{\hat{i}}$ from CW-Net were shown as percentages for easier interpretation [gigerenzer1995improve].
To evaluate the usefulness of the explanations in naturalistic settings [endsley2017autonomous], we deployed CW-Net on a real AV using the Lab2Car wrapper [heim2024lab2car]. All tests were performed on a closed course or private lot with an experienced safety driver. Figure 2 shows the experimental setup inside the AV. We next detail three notable situations in which the explanations proved beneficial.
5.1 Unexpected stopping for nearby vehicles
Figure 3: Results: a. The CLOSE concept activated when the car got stuck next to parked vehicles. The driver initially thought the pick-up/drop-off area was the cause, but the explanation suggested that it was the nearby vehicles. When the driver engaged self-driving away from the park vehicles, activation of the CLOSE concept decreased and the AV started moving again, counter to driver’s initial mental model and consistent with the explanations. Across tests, CLOSE correlated with speed and the intercept accurately predicted this event. b. The ASV concept activated when the car stopped next to a traffic cone. The driver initially thought the cone was the cause, but the explanation suggested that the AV was hallucinating a stopped vehicle. When we removed the cone in a counterfactual test, the same phantom braking and concept activation occurred. Across tests, ASV correlated with reductions in speed when spiking above 0.5 probability. c. The BIKE concept failed to activate in our first round of tests with a cyclist, but the car always stopped safely for the cyclist. Responding to the explanation, the driver engaged self-driving from slower speeds in a second round of tests. Follow-up analyses revealed that the AV stopped for the cyclist due to backup safety mechanisms unrelated to CW-Net, which indicates that the driver’s increased level of caution was appropriate. These three scenarios validate the driver’s updated mental models in response to the explanations. a-c. Findings were confirmed in an online study which replicated the events from the road. After seeing the explanations, subjects showed less agreement with the driver’s initial mental model, compared to a control group. Standard error of the mean and 3-second rolling averages is shown in relevant plots.
We observed that the AV repeatedly came to a stop shortly before a pedestrian pickup/drop-off zone (Figure 3a). The driver’s intuition was that the car stopped because of the pickup/drop-off zone, but the explanations indicated that the planner stopped because it detected that it was “Close to another vehicle” (the CLOSE concept). To test this hypothesis, the driver manually moved the car farther from the parked cars. At this point, the probability of CLOSE decreased and the AV began moving again, thus confirming the alternative hypothesis. A full timeline of events is detailed in Figure 3a. We fitted the intercept of the CLOSE probability against the speed of the AV globally and found it accurately predicts stopping and starting for this event. [^1] This situation used a variant of our architecture detailed in Figure S1.
5.2 Hallucinating a stopped vehicle ahead
At another location, the AV would reliably come to a stop next to a traffic cone (Figure 3b). The driver’s initial mental model was that the cone was responsible for the phantom brake. However, the “Approaching stopped vehicle” (ASV) concept peaked shortly before the car stopped. This suggested an alternative hypothesis, that the planner matched the current situation with training scenarios labeled ASV, which in turn promotes stopping behavior associated with such scenarios. As a counterfactual test, the cone was removed. The AV exhibited the same stopping behavior at the same location, along with similar ASV probability and speed profiles ($L_{2}$ similarities of 1.12 and 1.6 between the respective time-warped profiles, compared to an average $L_{2}>200$ for random events), thus confirming the alternative hypothesis. Note that although there was no vehicle in front of the AV, the explanation is causally faithful to the underlying motion planner and explains why it stopped (namely, because it incorrectly detected a stopped vehicle). Figure 3b illustrates a global analysis of ASV, showing it to be a powerful predictor of braking.
5.3 Reacting safely to cyclist
Finally, we tested the ability of the AV to stop safely for cyclists (Figure 3c). For each test, the driver engaged self-driving mode while approaching a cyclist. The driver was instructed to engage self-driving from a speed at which they felt safe, since this determines the subsequent speed of the AV. During the initial tests, the AV reliably stopped for the cyclist. However, the BIKE concept maintained a low probability throughout each test ($<$ 1%), indicating that CW-Net was failing to detect the cyclist. Over time, the driver became aware of the concept reading and gradually increased their caution by initiating self-driving from slower speeds. A post-hoc analysis revealed that, although the perception system detected the cyclist, the motion planner failed to take it into account due to a lack of appropriate input features for cyclists. As a result, it chose unsafe trajectories which would have collided with the cyclist. In reality, it was the built-in rule-based systems of the AV that overruled the motion planner and forced the AV to stop. This indicates that the increased caution dictated by the driver’s updated mental model was warranted.
6 Confirmation study
We conducted an online study (N=120) to confirm the statistical significance of our results. Following standard procedures in deployment-based research [FAA2011FlightTest, schneider2023don], we simulated the sequence of key events from the three situations described previously. We designed a between-subjects study to assess the effect of explanations on the mental model of subjects in the driver’s position. For each situation, subjects in the experimental group received the CW-Net explanation, while subjects in the control group received a generic explanation to balance cognitive load (see Figure S4 and Section User Study for details). Subjects then rated to what extent they agreed with the driver’s incorrect initial mental model. Across all situations, the experimental group gave significantly lower ratings than the control group (Figure 3, right column; mean result: 3.37$\pm$1.63 v. 5.46$\pm$0.89; t-test $p$ $<$ 0.001; Cohen’s d=1.58), indicating that the explanations support mental model alignment across the population.
7 Conclusion
Our work shows, for the first time, how explainable deep learning can provide useful explanations for the decisions of self-driving cars in the real world. CW-Net achieves this by grounding the reasoning of a pretrained black-box motion planner in human-interpretable concepts corresponding to types of scenario. By revealing otherwise inaccessible information about the decision-making process of the motion planner in real time, CW-Net helps align the mental model of the human driver with the machine driver. This allows the human driver to better anticipate and account for mistakes of the AV, ultimately resulting in safer driving. Mental model alignment could additionally build trust and understanding with passengers of driverless AVs by helping them anticipate the AV’s decisions. This could be particularly beneficial when AV behavior deviates from typical human behavior, such as when driving conservatively or getting stuck. Additionally, the explanations provided by CW-Net can help test engineers provide more precise feedback to the research scientists and engineers working on model improvement. This would be especially relevant for experimental motion planners that are still under development, such as the one used in our study.
CW-Net extends the original PW-Net [kenny2023towards], which reasons over specific scenario prototypes, to general scenario types. In addition to increasing robustness, reasoning over types has the added benefit of highlighting which parts of the training distribution influence behavior at each point in time. This information can be used by researchers for model improvement. For example, the inability of CW-Net to detect the BIKE concept suggests there may not be enough training scenarios with cyclists, or that the focal loss used was ineffective [lin2017focal], leading to poor performance around cyclists. At the same time, relying on types forgoes some of the benefits of using prototypes. For example, CW-Net can explain that it is stopping because the current scene is similar to ASV scenarios in the training data, but it cannot explain why it believes so  [rudin2019stop, chen2019looks]. This suggests a promising avenue of future research: much like humans, who rely on both exemplar-based and rule-based reasoning [smith1994similarity, davis2012learning], machines that reason over both prototypes and types could combine the strengths of both approaches, enhancing interpretability while maintaining flexibility in decision making.
Many safety-critical systems involving human-robot interaction require real-time explanations, including AI wingmen, drone navigation systems, and robotic surgeons. Similarly to AVs, many of these applications increasingly rely on deep learning and have a long tail of failure cases, with potentially catastrophic outcomes. This creates an ethical imperative to better understand how they work, so the humans in the loop can intervene when necessary. The success of CW-Net suggests that explainable deep learning may prove essential for meeting the safety and regulatory standards for deploying sophisticated safety-critical agents in the real world.
Acknowledgements
We thank Tom Bewley of JPMC AI research, Mycal Tucker of Anthropic, and Winthrop Gillis of Harvard for reviewing an early version of the manuscript.
Author contributions
E.M.K contributed to conceptualization of the research, model training, experimental evaluation, writing. A.D., S.U.L., T.P.M., S.R., Y.H., and M.S.T. all contributed to the technical implementation of the algorithm in the self-driving car. L.M. contributed to project organization, and writing. M.S.T. and J.A.S. contributed to conceptualization of the research, project organization, and writing.
Competing interests
The authors declare no competing interests.
Additional information
Supplementary information available at https://drive.google.com/drive/folders/1Lz6OGGi2gFeBOnC3ddyzFJMztqUTC_Am?usp=sharing and https://github.com/EoinKenny/CW_Net. Correspondence and requests for materials should be addressed to Eoin M. Kenny.
Supplementary Material
Figure S1: Parallel architecture: This is identical to the black-box planner (Figure 1b), except the scene-trajectory embeddings are fed to the concept classifier $C$ in parallel to the (original) reward model $R$. These concept classifications are then converted to probabilities (x100 to convert to percentages) and presented to the user.
Scenario Metric Our Model Original
nominal lane follow (968) Average L2 error 4.202058 4.200189
Average L2 error 3s 0.401251 0.400440
Average L2 error 5s 0.780286 0.778495
Average L2 error 10s 2.141165 2.136004
Progress along expert route 0.939595 0.939382
No at fault collisions 0.909091 0.909091
decelerating from high speed scenarios (1099) Average L2 error 3.139028 3.144618
Average L2 error 3s 0.533925 0.533905
Average L2 error 5s 0.930384 0.929786
Average L2 error 10s 2.056039 2.053457
Progress along expert route 0.990187 0.990213
No at fault collisions 0.945405 0.942675
Deceleration time difference 0.443201 0.439381
start accelerating from stationary scenarios (919) Average L2 error 8.919248 9.019697
Average L2 error 3s 0.426693 0.425365
Average L2 error 5s 1.244357 1.249598
Average L2 error 10s 4.611888 4.673006
Progress along expert route 0.740881 0.736437
No at fault collisions 0.807399 0.803047
Start from stationary max speed difference 0.239367 0.241552
Start from stationary time delay 3.477263 3.494675
Table S1: Full nuPlan results which compares the performance of our main model used in testing against the original black-box IRL agent.
Accuracy Precision Recall F1 Score
Steering 0.62 (cross entropy)
Speed 0.83 (cross entropy)
Approaching stopped vechicle 0.55 0.2 0.89 0.33
Intersection 0.51 0.42 0.61 0.49
CLOSE to another vechicle 0.45 0.1 0.81 0.17
Ranker Accuracy 0.94
Table S2: Concept performance for Dataset 1 for CW-Net
Accuracy Precision Recall F1 Score
Slow 0.83 0.73 0.93 0.82
Stopped 0.48 0.16 0.81 0.27
Fast 0.68 0.49 0.86 0.63
Stop sign 0.43 0.03 0.84 0.05
Traffic light 0.39 0.16 0.62 0.25
Intersection 0.42 0.20 0.65 0.31
Pedestrian 0.44 0.03 0.84 0.07
Following 0.43 0.02 0.84 0.04
BIKE 0.21 0.00 0.42 0.00
PUDO 0.58 0.30 0.86 0.44
Ranker Accuracy 0.95
Table S3: Concept performance for dataset 1 with CW-Net.
Figure S2: A comparison of the concept predictions for the same model deployed in a large test track with wide open, long straight roads, and a smaller parking lot with many turns.
Figure S3: Statistical results for user study: (left) Each individual question.(right). Averaged across all questions
Figure S4: Survey questions for the experimental and control group.