### v2

[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
Explainable deep learning improves human mental models of self-driving cars
[…]
Summary
Self-driving cars increasingly rely on deep neural networks to achieve human-like driving [1, 2, 3]. The opacity of such black-box planners makes it challenging for the human behind the wheel to accurately anticipate when they will fail [4, 5, 6], with potentially catastrophic consequences [7, 8, 9]. While research into interpreting these systems has surged, most of it is confined to simulations or toy setups due to the difficulty of real-world deployment [10, 11], leaving the practical utility of such techniques unknown. Here, we introduce the Concept-Wrapper Network (CW-Net), a method for explaining the behavior of machine-learning-based planners by grounding their reasoning in human-interpretable concepts. We deploy CW-Net on a real self-driving car and show that the resulting explanations improve the human driver’s mental model of the car, allowing them to better predict its behavior. To our knowledge, this is the first demonstration that explainable deep learning integrated into self-driving cars can be both understandable and useful in a realistic deployment setting. CW-Net accomplishes this level of intelligibility while providing explanations which are causally faithful and do not sacrifice driving performance. Overall, our study establishes a general pathway to interpretability for autonomous agents by way of concept-based explanations, which could help make them more transparent and safe.
keywords:
Self-driving cars, Mental models, Situational awareness, Interpretable ML
1 Introduction
There are hundreds of companies developing autonomous vehicle (AV) technology globally [12], promising to revolutionize transportation for everyone. The industry currently spans two segments: driverless ride-hail systems and consumer vehicles [13]. In consumer vehicles, machine learning (ML) solutions have markedly improved advanced driver-assistance systems, yet they still require human intervention in unusual or challenging situations where learned planners may not determine the correct action [14]. As driving is safety-critical, such infrequent failures matter, making it essential that the human driver is able to anticipate and be prepared for such situations [15]. However, the opaque nature of ML planners makes it challenging to interpret and communicate the causes of their decisions, hampering the ability of human drivers to understand and predict AV behavior while achieving real-time situational awareness [16, 17, 18].
Lack of effective communication between the AV and the human driver has contributed to multiple high-profile incidents, some resulting in fatalities [7, 8, 9], highlighting the urgent need to make ML planners interpretable [11]. Previous studies have sought to address this using surveys and simulated scenarios [19, 20, 21, 22, 23, 24, 25, 26], a human driver emulating the AV [27, 28], or large language models providing post-hoc explanations [3]. However, these studies were theoretical, did not provide causally faithful explanations, were only evaluated in simulation, or did not convincingly show the practical utility of the explanations. This leaves open the question of how to provide explanations that are understandable, useful, and faithful to the decision-making process of the AV in a realistic setting.
To answer this question, we scale up our recent work on interpretable-by-design deep reinforcement learning [10] using motifs from the literature on concept-bottleneck models [29] to propose the Concept-Wrapper Network (CW-Net). CW-Net grounds the reasoning of a black-box ML planner in human-interpretable concepts, such as “Approaching stopped vehicle” or “Close to cyclist”. This method is rooted in case-based reasoning, a classical artificial intelligence (AI) approach [30, 31, 32] inspired by cognitive models of human reasoning and memory [33]. CW-Net can be applied to arbitrary pretrained deep neural networks, does not require retraining from scratch, and does not degrade the performance of the original black-box ML planner. Since the inferred concepts are the sole input to the final decision-making module of CW-Net that directly dictates AV behavior, we refer to them as causally faithful explanations.
We apply CW-Net to a ML planner trained to imitate human driving behavior using inverse reinforcement learning [1, 2]. We replace the final (reward) layer of the pretrained deep neural network with a concept classifier, followed by a new reward layer. We then jointly train the classifier and the new reward layer to predict scenario types and driving decisions, respectively, without modifying the rest of the network. Evaluation on a large-scale benchmark [10] confirms that CW-Net is able to classify concepts without compromising driving behavior. To study the utility of the explanations, we deploy CW-Net on a real self-driving car with a safety driver in a semi-naturalistic study [35]. We demonstrate three situations in which the driver’s initially inaccurate mental model is subsequently improved by the CW-Net explanations, which in turn allows the driver to more accurately predict AV behavior. We then observe similar mental model improvement in a larger online study simulating these scenarios. Lastly, we link these results to more complex real-world scenarios by deploying CW-Net on public roads in Las Vegas. We use these new data in a large online study ($N=100$) to demonstrate that improvements in mental model goodness lead to improved predictive ability and situational awareness. Overall, our work demonstrates how explainable AI can help users of an advanced autonomous system improve their mental model of the system and better anticipate its future behavior.
2 Black-box planner
Figure 1: Planner architecture, adapted with permission from Tomov et al. (2025) [1]. a. Autonomous vehicle stack. Sensory input is processed by the perception module to generate scene context $s$. The planning module processes $s$ to compute trajectory $\hat{\tau}$, which is followed by the control module. b. Black-box ML planner. Here, $s$ is fed to trajectory generator $G$, which produces candidate trajectories $\{\tau_{1}\dots\tau_{k}\}$, and scene encoder $H$, which produces a scene embedding $\mathbf{h}$. These are fed into encoder $E$, which produces scene-trajectory embeddings $\mathbf{z}_{i}$ for each $\tau_{i}$, which are in turn fed into the reward layer $R$. $R$ computes a reward $r_{i}$ for each trajectory. The model is trained to output higher rewards for trajectories closer to the the ground-truth human trajectories. c. CW-Net. Identical to b., except $\mathbf{z}_{i}$ is fed to a concept classifier $C$, which produces concept assignments $\mathbf{c}_{i}$. These are fed to a new reward layer $R^{\prime}$ to produce rewards $r^{\prime}_{i}$. In parallel, $\mathbf{c}_{i}$ is processed to generate the explanations. The reward layer is trained to prefer the same trajectories as the black-box ML planner, while the concept layer is supervised with ground-truth scenario labels. The weights of the faded components are frozen during training. CLOSE, “Close to another vehicle”. ASV, “Approaching stopped vehicle”. BIKE, “Close to cyclist”.
We focus on the planning module of the AV stack (Figure 1a), which takes as input a scene context $s$ and outputs a trajectory $\hat{\tau}$. Here, $s$ is a symbolic object-oriented representation of the scene, computed by the perception module, while $\hat{\tau}$ is the trajectory that the subsequent controller module should follow. We use a deep neural network architecture [1, 2] consisting of a scene encoder $H(s)\rightarrow\mathbf{h}$ and a trajectory generator $G(s)\rightarrow\{\tau_{1}\dots\tau_{k}\}$, followed by a scene-trajectory encoder $E(\mathbf{h},\tau_{i})\rightarrow\mathbf{z}_{i}$ and a final reward layer $R(\mathbf{z}_{i})\rightarrow r_{i}$ (Figure 1b). $H$ computes a scene embedding $\mathbf{h}$ and $G$ computes a set of $k$ candidate trajectories $\{\tau_{1}\dots\tau_{k}\}$. Those are combined in $E$ to compute an embedding $\mathbf{z}_{i}$ for each trajectory $\tau_{i}$. Finally, $R$ computes an estimated reward $r_{i}$ for each trajectory $\tau_{i}$, quantifying how human-like it is. In other words, $r_{i}$ is higher when $\tau_{i}$ is more similar to how a human would drive in this situation.
During inference, the trajectory with highest reward is selected on each iteration:
$\hat{\tau}=\underset{\tau\in{\{\tau_{1}...\tau_{k}}\}}{\arg\max}R(E(\mathbf{h},\tau))$
We use inverse reinforcement learning to train the planner on 80 hours of human expert driving [1, 2].
3 Planning over human-friendly concepts
To make the ML planner interpretable, we modify the planner architecture to additionally provide explanations for its behavior [10]. Our working definition of explainable AI follows Gunning & Aha [36], who define it as ‘AI systems that can explain their rationale to a human user, characterize their strengths and weaknesses, and convey an understanding of how they will behave in the future.’
Specifically, we replace $R$ with a concept classifier $C(\mathbf{z}_{i})\rightarrow\mathbf{c}_{i}$, followed by a new reward layer $R^{\prime}(\mathbf{c}_{i})\rightarrow r^{\prime}_{i}$ (Figure 1c). The concept classifier $C$ computes a logit vector $\mathbf{c}_{i}$ that is passed through a softmax and/or sigmoid layer that, in turn, assigns probabilities to different human-interpretable concepts. Since $R^{\prime}$ computes trajectory rewards from $\mathbf{c}_{i}$, the final decisions are based solely on these concept assignments and hence they constitute a causally faithful explanation. The rest of the network remains the same.
Similarly to the black-box planner, trajectories are selected according to:
$\hat{\tau}=\underset{\tau\in{\{\tau_{1}...\tau_{k}}\}}{\arg\max}R^{\prime}(C(E(\mathbf{h},\tau)))$
We train CW-Net to jointly predict concept labels and mimic the driving decisions of the black-box planner. Specifically, $\mathbf{c}_{i}$ is supervised with labels corresponding to types of scenarios, such as “Approaching stopped vehicle” or “Close to cyclist” (see Section 8.5 for a full list of concepts). This ensures that CW-Net assigns a unique interpretable concept to each unit in $\mathbf{c}_{i}$. At the same time, $r^{\prime}_{i}$ is supervised with trajectories selected by the black-box planner using a cross-entropy loss. During training, the rest of the deep neural network ($H$, $G$, and $E$) is kept frozen (see Section 8.2 for more details). We focus our study on this particular ML planner architecture, noting that prior work indicates CW-Net would generalize effectively to other architectures [10].
4 Classifying concepts while preserving planner performance
As a baseline, we first evaluated the black-box planner (without CW-Net) using closed-loop simulations with the nuPlan simulator on the nuPlan dataset [10] (Table S1). Overall, the results were competitive with the top submissions to the nuPlan challenge, although performance was slightly lacking when starting from a stop (see Section 8). This suggests that there is room for improvement and, importantly, opportunities to study explanations of undesirable behavior.
We then evaluated the driving performance of CW-Net wrapped around the black-box planner (Table S1). The results were equivalent, with less than 1% difference across all metrics, confirming that our method did not degrade driving performance. We also evaluated concept classification on held-out datasets (Tables S2S3). Mean accuracy was 54%, with 23% precision, 77% recall, and an F1 score of 0.31 (see Section 8.3). Overall, these results indicate that CW-Net can be used to ground the decision making of high-performance ML planners in human-interpretable concepts, without sacrificing driving performance.
5 Mental model improvement in surprising real-world situations
Our central hypothesis is that the explanations from CW-Net would improve the human driver’s mental model of the AV and, by extension, their situational awareness. This would be particularly salient in surprising situations, which is when explanations are most useful [37]. Specifically, the explanations should improve the driver’s ability to understand and address the reasons for AV failures, and also to predict its future behavior.
To test this hypothesis, we deploy CW-Net on a real AV on a private track using the Lab2Car wrapper [35] and observe how safety drivers react to surprising events and the corresponding CW-Net explanations (Figure 2). These situations were not planned, but instead occurred naturally, with minimal intervention from the researchers, who only dictated high-level plans for each day. This semi-naturalistic study allowed us to assess the utility of the explanations in naturally occurring surprising situations. To measure AV predictability, we record the drivers’ ability to make counterfactual predictions about AV behavior in these situations. To note the drivers’ mental models, together with their predictive ability, we additionally record their think-aloud thoughts before and after considering CW-Net explanations [11]. We focus on concepts that relate to other road users and can be easily tested counterfactually (see Section 8.5).
Figure 2: Deployment setup. A safety driver, a support engineer, and a researcher were present. The safety driver drove the AV manually between road tests, engaged self-driving mode at the start of each test, monitored AV performance during the test, and took over in case of unsafe driving. The support engineer deployed CW-Net and set scenario destinations. The researcher directed testing. The dashboard included a map with overlaid object detections ($s$) from the perception module and the output trajectory ($\hat{\tau}$). Explanations $\mathbf{c}_{\hat{i}}$ from CW-Net were shown as percentages for easier interpretation [39].
5.1 Unexpected stopping for nearby vehicles
Figure 3: Results: a. The CLOSE concept activated when the car got stuck next to parked vehicles. The driver initially thought the pickup/drop-off area was the cause, but the explanation suggested that it was the nearby vehicles. When the driver engaged self-driving away from the park vehicles, activation of the CLOSE concept decreased and the AV started moving again, counter to driver’s initial mental model and consistent with the explanations. Across tests, CLOSE correlated with speed and the intercept accurately predicted this event. b. The ASV concept activated when the car stopped next to a traffic cone. The driver initially thought the cone was the cause, but the explanation suggested that the AV was hallucinating a stopped vehicle. When we removed the cone in a counterfactual test, the same phantom braking and concept activation occurred. Across tests, ASV correlated with reductions in speed when spiking above 0.5 probability. c. The BIKE concept failed to activate in our first round of tests with a cyclist, but the car always stopped safely for the cyclist. After observing the explanation, the driver engaged self-driving from slower speeds in a second round of tests. Follow-up analyses revealed that the AV stopped for the cyclist due to backup safety mechanisms unrelated to CW-Net, which indicates that the driver’s increased level of caution was appropriate. a-c. (Right column) Naturalistic scenarios from public-road deployment, analogous to the scenarios observed on the private track. We used the PEDESTRIAN concept instead of the BIKE concept as it behaved similarly and pedestrians were more frequently encountered on public roads. Likewise, we used pedestrians in place of the traffic cone in our ASV tests as there were no cones encountered on the public roads. Standard error of the mean and 3-second rolling averages are shown in relevant plots.
We observed that the AV repeatedly came to a stop shortly before a pedestrian pickup/drop-off zone (Figure 3a). The driver’s intuition was that the car stopped because of the pickup/drop-off zone, but the explanations indicated that the planner stopped because it detected that it was “Close to another vehicle” (the CLOSE concept). To test this hypothesis, the driver manually moved the car farther from the parked cars. At this point, the probability of CLOSE decreased and the AV began moving again, thus supporting the alternative hypothesis. A full timeline of events is detailed in Figure 3a. We fitted the intercept of the CLOSE probability against the speed of the AV globally and found it accurately predicts stopping and starting for this event.
5.2 Hallucinating a stopped vehicle ahead
At another location, the AV would reliably come to a stop next to a traffic cone (Figure 3b). The driver’s initial mental model was that the cone was responsible for the phantom brake. However, the “Approaching stopped vehicle” (ASV) concept peaked shortly before the car stopped. This suggested an alternative hypothesis, that the planner matched the current situation with training scenarios labeled ASV, which in turn promotes stopping behavior associated with such scenarios. As a counterfactual test, the cone was removed. The AV exhibited the same stopping behavior at the same location, along with similar ASV probability and speed profiles ($L_{2}$ similarities of 1.12 and 1.6 between the respective time-warped profiles, compared to an average $L_{2}>200$ for random events), thus supporting the alternative hypothesis. Note that although there was no vehicle in front of the AV, the explanation is causally faithful to the underlying planner and explains why it stopped (namely, because it incorrectly detected a stopped vehicle). Figure 3b illustrates a global analysis of ASV, showing it to be a powerful predictor of braking.
5.3 Reacting safely to cyclist
Finally, we tested the ability of the AV to stop safely for cyclists (Figure 3c). For each test, the driver engaged self-driving mode while approaching a cyclist. The driver was instructed to engage self-driving from a speed at which they felt safe, since this determines the subsequent speed of the AV. During the initial tests, the AV reliably stopped for the cyclist. However, the BIKE concept maintained a low probability throughout each test (<< 1%), indicating that CW-Net was failing to detect the cyclist. Over time, the driver became aware of the concept reading and gradually increased their caution by initiating self-driving from slower speeds. A post-hoc analysis revealed that, although the perception system detected the cyclist, the ML planner was not configured to consume inputs for cyclists. As a result, it chose unsafe trajectories which would have collided with the cyclist. In reality, the AV stopped due to a built-in safety backup system, which commands a brake if collision is imminent. This indicates that the increased caution dictated by the driver’s updated mental model was warranted.
6 Mental model improvement in simulation studies
The semi-naturalistic study above illustrates how CW-Net explanations can improve the human driver’s mental model of the AV in surprising real-world situations, leading to better understanding and predictability of AV behavior. To validate that these results replicate in larger populations and generalize to naturalistic scenarios, we conduct several larger follow-up studies (Figures S2S3). First, we show that in simulations of the real-world scenarios in Section 5, CW-Net explanations consistently improve mental models and predictions for both experts and non-experts alike. Second, we collect data from CW-Net in naturalistic real-world scenarios on public roads and show that the explanations improve situational awareness – an indirect measure of mental model goodness [12] – in surprising situations, without degrading it in unsurprising situations.
6.1 Explanations improve mental models and predictability
The purpose of the first study is to (1) demonstrate that the effects observed on the road reproduce in larger populations, and (2) establish a link between direct measures of mental model goodness and performance on downstream tasks, such as predicting AV behavior. We conducted an online survey in which we showed participants replays of the scenarios from the private track (front camera recordings with overlaid CW-Net explanations; Figure S11).
For each scenario, we asked participants to choose the reason for AV behavior that best aligns with their current beliefs (i.e., “Why did the AV do that?”) and to make counterfactual predictions (i.e., “What would the AV do if …?”). Following Hoffman et al. [11], we refer to the former as the nearest-neighbor task – since it forces participants to pick the nearest explanation to their beliefs – and the latter as the prediction task. Both tasks consist of a multiple-choice question followed by a confidence score. The prediction task additionally includes a free-form text response where participants can describe the reasons for their prediction. The nearest-neighbor task and the free-form text response directly probe the participants’ mental models, while the multiple-choice prediction question serves as a proxy performance metric for mental model goodness. Critically, we collected responses before and after observing the explanations. This within-subject design mirrors the experience of drivers during the real-world tests (Figures S2, S3). The study was conducted with experts ($N=9$ Motional drivers/engineers) and non-experts ($N=30$ pseudo-random participants from Prolific.com).
On the nearest-neighbor task, CW-Net explanations improved mental models for almost all participants (8/9 experts and 27/30 non-experts; Figure S4). Free-form text responses were initially similar to the initial beliefs of the safety drivers during the on-road tests ($p<10^{-9}$) and then shifted towards their final beliefs ($p<0.0002$) and the ground-truth reasons for AV behavior ($p<10^{-5}$; exact binomial tests for non-experts; Figures S6S7). The distribution of mental model updates on both tasks was similar across both groups ($\beta=0.04,p=0.8$, ordinary least-squares regression; Figure S8). Importantly, mental model improvement on the nearest-neighbor task correlated with improvements in prediction accuracy ($\beta=2.02\pm 0.87,p=0.02$ for experts, $\beta=9.86\pm 2.07,p<0.001$ for non-experts; linear mixed-effects models [LMEs]; Figure S9, left). A similar effect was observed for the free-form text responses ($\beta=1.70\pm 0.91,p=0.06$ for experts, $\beta=5.03\pm 1.23,p<0.001$ for non-experts; LMEs; Figure S9, right). Together, these results support the conclusion that CW-Net explanations can reliably improve mental models of the AV and that prediction performance can serve as a reliable proxy for mental model goodness.
6.2 Explanations improve situational awareness in naturalistic scenarios
To verify that the effects observed on the private track generalize to more complex, naturalistic scenarios, we deployed CW-Net on public roads in Las Vegas (Figures S3; Figure 3, right). Due to safety reasons and the experimental nature of the ML planner, the safety driver operated the AV in ‘manual’ mode for several hours, with CW-Net running in the background. This allowed us to collect naturalistic scenarios analogous to those encountered on the private track [41, 28] (Figure S10). We use replays of those scenarios to conduct a large-scale online study ($N=100$) within the Situation Awareness Global Assessment Technique (SAGAT) framework tailored for explainable AI [12, 42].
The SAGAT framework is the gold standard for evaluating situational awareness in complex, dynamic tasks [12, 42]. It measures the situational awareness of a human operator by freezing a scenario, querying the operator about their perception (i.e., “What are the inputs to the AV?”), comprehension (i.e., “Why is the AV doing that?”), and projection (i.e., “What would the AV do if …?”) of the situation (Figure S11) and comparing their responses with the ground truth. As counterfactual predictive performance in Section 6.1 aligns with mental model goodness, we use projection in SAGAT as a proxy for mental model elicitation and quality [11]. For each type of scenario and corresponding concept, we sample two instances in which AV behavior was surprising (e.g., stopping unnecessarily) and two corresponding instances in which it was unsurprising (Table S6). We use a between-subject design in which the experimental group received explanations from CW-Net, while the control group received placeholder explanations of AV kinematics to balance cognitive load [14] (Figures S12-S14).
We found that the CW-Net explanations significantly improved measures of situational awareness for surprising events (Figure S15, left; Table S7, top), with large effect sizes for perception (Cohen’s d = 1.290), comprehension (Cohen’s d = 0.996), and a medium effect size for projection (Cohen’s d = 0.606). Conversely, explanations did not impact situational awareness for unsurprising events (Figure S15, right; Table S7, bottom). Together, these results show that the explanations from CW-Net are robust across various conditions and can reliably improve mental models in surprising situations, without significant negative effects in unsurprising situations. Moreover, they demonstrate pragmatic usage of the improved mental models.
7 Conclusion
Our work shows, for the first time, how explainable deep learning can provide useful explanations for AVs in a real-world setting. CW-Net achieves this by grounding the reasoning of a pretrained black-box ML planner in human-interpretable concepts that are directly used to make driving decisions. By revealing otherwise inaccessible information about the decision-making process of the AV in real time, CW-Net helps improve the human driver’s mental model of the AV. This, in turn, improves the driver’s situational awareness and reveals limitations of the robotic system, helping the driver better anticipate its mistakes.
Many systems involving human-robot interaction require real-time explanations, and many of these applications increasingly rely on deep learning, with a long tail of potentially catastrophic failure cases. Indeed, many regulatory bodies have already made explainable AI a core component of their legislation, with AVs likely to follow suit as they are widely deployed with various users [11]. As such, the success of CW-Net suggests that similar algorithms may prove essential for meeting the regulatory standards for deploying AVs, while building appropriate trust in the technology. In future work, it would be prudent to extend CW-Net to cover a larger set of concepts – perhaps in an unsupervised manner to overcome the challenges of labeling – to better cover the vast array of concepts relevant to AV settings.
Acknowledgements
We thank Tom Bewley of JPMC AI research, Mycal Tucker of Anthropic, and Winthrop Gillis of Harvard for reviewing an early version of the manuscript.
Author contributions
E.M.K contributed to conceptualization of the research, model training, experimental evaluation, writing. A.D., S.U.L., T.P.M., S.R., Y.H., and M.S.T. all contributed to the technical implementation of the algorithm in the self-driving car. L.M. contributed to project organization, and writing. M.S.T. and J.A.S. contributed to conceptualization of the research, project organization, and writing.
Competing interests
The authors declare no competing interests.
Additional information
Supplementary information available at https://drive.google.com/drive/folders/1Lz6OGGi2gFeBOnC3ddyzFJMztqUTC_Am?usp=sharing and https://github.com/EoinKenny/CW_Net. Correspondence and requests for materials should be addressed to Eoin M. Kenny.
Supplementary Material
Figure S1: Parallel architecture: This is identical to the black-box planner (Figure 1b; greyed out here), except the scene-trajectory embeddings are fed to the concept classifier $C$ in parallel to the (original) reward model $R$. The black-box part of the architecture (greyed out) is kept frozen. These concept classifications are then converted to probabilities (x100 to convert to percentages) and presented to the user.
Scenario type Metric CW-Net ML planner
Nominal lane follow (968 scenarios) Average L2 error 4.202058 4.200189
Average L2 error 3s 0.401251 0.400440
Average L2 error 5s 0.780286 0.778495
Average L2 error 10s 2.141165 2.136004
Progress along expert route 0.939595 0.939382
No at fault collisions 0.909091 0.909091
Decelerating from high speed (1099 scenarios) Average L2 error 3.139028 3.144618
Average L2 error 3s 0.533925 0.533905
Average L2 error 5s 0.930384 0.929786
Average L2 error 10s 2.056039 2.053457
Progress along expert route 0.990187 0.990213
No at fault collisions 0.945405 0.942675
Deceleration time difference 0.443201 0.439381
Start accelerating from stationary (919 scenarios) Average L2 error 8.919248 9.019697
Average L2 error 3s 0.426693 0.425365
Average L2 error 5s 1.244357 1.249598
Average L2 error 10s 4.611888 4.673006
Progress along expert route 0.740881 0.736437
No at fault collisions 0.807399 0.803047
Start from stationary max speed difference 0.239367 0.241552
Start from stationary time delay 3.477263 3.494675
Table S1: Full nuPlan results comparing driving performance of our CW-Net planner to the original black-box ML planner.
Accuracy Precision Recall F1 Score
Steering 0.62 (cross entropy)
Speed 0.83 (cross entropy)
Approaching stopped vechicle (ASV) 0.55 0.20 0.89 0.33
Intersection 0.51 0.42 0.61 0.49
CLOSE to another vechicle 0.45 0.10 0.81 0.17
Ranker accuracy 0.94
Table S2: CW-Net concept classification performance on Dataset 1
Accuracy Precision Recall F1 Score
Slow 0.83 0.73 0.93 0.82
Stopped 0.48 0.16 0.81 0.27
Fast 0.68 0.49 0.86 0.63
Stop sign 0.43 0.03 0.84 0.05
Traffic light 0.39 0.16 0.62 0.25
Intersection 0.42 0.20 0.65 0.31
Pedestrian 0.44 0.03 0.84 0.07
Following 0.43 0.02 0.84 0.04
BIKE 0.21 0.00 0.42 0.00
PUDO 0.58 0.30 0.86 0.44
Ranker accuracy 0.95
Table S3: CW-Net concept classification performance on Dataset 2.
Figure S2: Study methodology. First we collected a series of surprising situations expert safety drivers encountered during their day-to-day job testing the AV where the explanations proved to be useful to understand why the AV made certain decisions and how it would behave in alternative scenarios. During these situations, we noted the driver’s mental model through their think-aloud thoughts and behaviour. A mental model elicitation study was run with more expert drivers and non-experts online to more robustly verify these findings and if mental model goodness did indeed correlate with counterfactual predictive performance. Then, in a large-scale SAGAT study, we used projection as a surrogate metric for mental model quality, since it perfectly matches the prior prediction tasks.
Figure S3: Study pipeline example. We noted the safety drivers’ think-aloud thoughts and predictions in surprising situations encountered during the on-road tests. In the subsequent online mental model elicitation study, we used their beliefs about AV behaviour before/after the explanation in the nearest-neighbor task and their predictions about future AV future behavior in the prediction task. The nearest-neighbor task resembles the comprehension task in the follow-up SAGAT study. The prediction task corresponds exactly to the projection task in the SAGAT study.
Scenario/concept Before explanation After explanation
CLOSE 0.00% 11.11%
ASV 33.33% 44.44%
BIKE 77.78% 77.78%
Overall average 37.04% 44.44%
Table S4: Performance improvement on prediction task ($N=9$ experts).
Scenario/concept Before explanation After explanation
CLOSE 3.33% 33.33%
ASV 36.67% 60.00%
BIKE 40.00% 66.67%
Overall average 26.67% 53.33%
Table S5: Performance improvement on prediction task ($N=30$ non-experts).
Figure S4: Mental model improvement on nearest-neighbor task. Visualizing belief updates in response to the explanations for experts (left) and non-experts (right) on the nearest-neighbor task. Arrows originate at the participants’ initial accuracy/confidence (before observing the explanation) and terminate at their final accuracy/confidence (after observing the explanation). For each participant, initial and final accuracy/confidence is averaged across all 3 scenarios. Green arrows indicate improvement in choice accuracy/confidence, while red arrows indicate worsening. Gray points represent participants with no change in beliefs.
Figure S5: Performance improvement on prediction task. Visualizing belief updates in response to the explanations for experts (left) and non-experts (right) on the prediction task. Notation as in Figure S4.
Figure S6: Mental model improvement on free-form text rationale ($N=9$ experts). Confusion matrices with participant beliefs (mental model and predictions) extracted from free-form responses in the mental model elicitation study study (y-axis) compared to beliefs from the safety drivers from the on-road tests and the ground truth explanation for AV behavior (x-axis). Numbers indicate raw response counts.
Figure S7: Mental model improvement on free-form text rationale ($N=30$ non-experts). Notation as in Figure S6.
Figure S8: Correlation between mental model metrics across groups. Heatmaps displaying the frequency of alignment between changes in nearest-neighbor (NN) task responses and the free-form text rationale. To determine if the relationship between these two mental model proxies differs by expertise, we performed ordinary least-squares (OLS) regression with clustered standard errors, including an interaction term. The interaction between group (experts vs. non-experts) and score change was not significant ($\beta=0.04,p=0.842$), indicating that the relationship between quantitative score improvements and qualitative rationale improvements is largely consistent across both groups.
Figure S9: Mental model improvement correlates with prediction improvement. Linear mixed effects (LME) models were used to analyze the relationship between changes in mental model proxies (nearest-neighbor task and free-form text rationale) and prediction accuracy/confidence. Specifically, the y-axis indicates each participant’s Likert confidence after explanation minus before, positive scores indicate they moved towards the correct prediction, and green dots indicate the prediction flipped correctly. For experts ($N=9$ participants x 3 scenarios), improvement in the nearest-neighbor (NN) task significantly correlates with prediction improvement ($\beta=2.02,SE=0.87,p=0.02$). The text rationale shows a similar positive effect ($\beta=1.70,SE=0.91,p=0.06$), although it does not reach significance. For non-experts ($N=30$ participants x 3 scenarios), both NN improvement ($\beta=9.86,SE=2.07,p<0.001$) and text rationale improvement ($\beta=5.03,SE=1.23,p<0.001$) are significantly correlated with prediction improvement. Positive betas indicate that improving the mental model proxy is associated with a higher increase in prediction accuracy/confidence.
Figure S10: Naturalistic scenario examples. The original surprising scenarios discovered on the private track and the analogous scenarios discovered on public roads.
Figure S11: SAGAT materials example. The two conditions (explanation vs. no explanation) were presented with the same video, followed by a “blackout”, and then 6 questions related to situational awareness.
Testing concept robustness
Concept activation AV speed Concept accuracy
Original
ASV High Brakes Bad
CLOSE High Frozen Good
BIKE Low Moving Low
Modified for robustness
ASV High Brakes Good
CLOSE Low Frozen Good
BIKE Low Stopped Low
Table S6: Selecting unsurprising events for SAGAT study. To select unsurprising events we modifed what was individual unusual about each surprising event which fell into three categories. ASV: This concept was surprising because of the misclassification of the concept, so we selected a scenario in which it accurately classified the concept. CLOSE: This was surprising because the concept activation was higher than usual while the AV stopped, so we chose scenarios in which the AV stopped, but the concept activation was closer to its mean value. BIKE: This was unusual because the AV stopped seemingly due to the cyclist, so we chose scenarios in which the AV again stopped for other reasons. In all three setups, we only altered the one respective dimension, the other two remained static.
Figure S12: User materials example for the CLOSE concept, during a surprising and unsurprising event.
Figure S13: User materials example for the ASV concept, during a surprising and unsurprising event.
Figure S14: User materials example for the PEDESTRIAN concept, during a surprising and unsurprising event.
Figure S15: Online SAGAT study results. Users in the experimental group reported significantly higher accuracy on perception, comprehension, and projection during surprising events. During unsurprising events, there were no significant differences. Although a negative trend was noted for comprehension in the experimental group, it did not reach significance after Bonferroni correction (see Table S7). Taken with the results of the mental model elicitation studies, which found counterfactual prediction accuracy correlated with mental model goodness, these SAGAT results strongly indicate the improvements in situational awareness were due to underlying improvements in user mental models of the AV.
Valence Dimension Effect Cohen’s d $p$-value Bonferroni
direction significant
Surprising events Perception Exp >> Control 1.290 (large) <0.001<0.001 Yes
Comprehension Exp >> Control 0.996 (large) <0.001<0.001 Yes
Projection Exp >> Control 0.606 (medium) 0.003 Yes
Unsurprising events Perception No effect 0.085 (negligible) 0.667 No
Comprehension Control >> Exp 0.514-0.514 (medium) 0.011 No
Projection No effect 0.142-0.142 (negligible) 0.481 No
Table S7: SAGAT study results. During the ‘surprising’ events, explanations had an all-around positive impact on people’s situational awareness and thus their mental models. Conversely, during ‘unsurprising’ events, there were no significant differences between groups, after Bonferroni correction.
FEATURE Model 1 Model 2 Model 3
STOPPED 0.0862 0.0014 0.0324
SLOW 0.1575 0.0014 0.0324
FAST 0.1903
STOP SIGN 0.2368
TRAFFIC LIGHT 0.0938
INTERSECTION 0.0856 0.0247 0.0535
PEDESTRIAN 0.0998
FOLLOWING 0.1056
BIKE 0.1222
PUDO 0.2347
LEFT 0.0472 0.0332
RIGHT 0.0408 0.0410
STRAIGHT 0.0428 0.0592
ASV 0.0269 0.1050
CLOSE 0.0659 0.0702
Table S8: Negligible concept distribution shift: Wasserstein distances between the distribution of concept probabilities during the semi-naturalistic tests on the private track and the subsequent public-road test. Overall, the differences varied from mostly insignificant to low/medium in STOP SIGN and PUDO.